Edwards Construction currently has debt outstanding with a market value of $88,000 and a cost of 9 percent. The company has EBIT of $7,920 that is expected to continue in perpetuity. Assume there are no taxes.

  

a-1.

What is the value of the company's equity? (Do not round intermediate calculations. Leave no cell blank - be certain to enter "0" wherever required.)

a-2. What is the debt-to-value ratio? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
b. What are the equity value and debt-to-value ratio if the company's growth rate is 4 percent? (Do not round intermediate calculations and round your "Debt-to-value" answer to 3 decimal places, e.g., 32.161.)
c. What are the equity value and debt-to-value ratio if the company's growth rate is 8 percent? (Do not round intermediate calculations and round your "Debt-to-value" answer to 3 decimal places, e.g., 32.161.)
Answer #1

a-1 Interest Payment =9%*88000 =7920
Cash flow to Equity holders =EBIT -Interest =7920-7920 =0
Hence Value of Equity =0

a-2 Debt To Value Ratio =Debt/(Debt+Equity)=Debt/Debt =1

b. With growth of 4% EBIT next year =7920*(1+4%) =8236.80
Value of equity =(8236.80-Interest)/(9%-4%) =(8236.80-7920)/5% =6336
Debt to total Assets =88000/(88000+6336)=0.933

c. With growth of 8% EBIT next year =7920*(1+8%) =8553.60
Value of equity =(8553.60-Interest)/(9%-8%) =(8553.60-7920)/1% =63360
Debt to total Assets =88000/(88000+63360)=0.581